OpenStudy (anonymous):

What is the value of the expression when a = 4? 2a^3 i think 216

OpenStudy (anonymous):

I think its 512

OpenStudy (anonymous):

ya 512

OpenStudy (anonymous):

2*4*4*4

OpenStudy (anonymous):

2*4^3=2*4*4*4=8*16=128

OpenStudy (anonymous):

Two times four is 8. to the power of 3 which mean 8*8*8=512

OpenStudy (anonymous):

No, that's not what the question says. That would be true if it was (2a)^3 but it is not.

OpenStudy (anonymous):

youre right tom982 except that you would have to multiply it and then cube root it so @KailaMorgan7777 is right

OpenStudy (sleepyjess):

a=4 4^3=4*4*4 4*4=16 16*4=64 64*2=

OpenStudy (anonymous):